OpenStudy (anonymous):

Log question: \[\log_2({x ^{2}-7x+14}) = 1 \]

myininaya (myininaya):

\[\log_2(x^2-7x+14)=\log_2(2)\]

myininaya (myininaya):

set insides equal

OpenStudy (anonymous):

2^1 = (x^2-7x+14) that's how I was taught to do it.

myininaya (myininaya):

that is fine also

OpenStudy (anonymous):

thankyou!